Looking for help solving the following question:

Question 28: Solve the following IVP showing the first six non-zero terms explicitly.

x – 2tx + x = 0, x(0) = 1, x(0) = -1
Answer #1

X - 2tx tx=0 x(0)=1 , X(0) = -1 anth n=o Let x= In X = n(n-1) anth-2 nan n=1 n=2 X - 2tx tx = 0 n=2 n! n=0 n-2 ne neo n+2-2az 2.1 - 1 ai (1+1)(1+2) a 2-1 2.3 - a , te 3 2.1-1 31 3 12 - aa a2 - m/ -Ġ 2.2-1 (2+1)(2+2) 2.3-1 (3+1)(3+2) 3 41 2-2-1 3 /